The balloonist was travelling at 62.6 m/s when the parachute opened. a) What was the average acceleration? b) Why was the value of the average less than that for gravity?
1
Expert's answer
2019-02-08T09:50:02-0500

a) The balloonist started from 0 m/s in vertical direction. Then in time

"\Delta t"

he reached the speed of 62.6 m/s, thus the acceleration was

"a=\frac{v_1-v_0}{\Delta t}=\frac{62.6-0}{\Delta t}=\frac{62.6}{\Delta t}."

b) From the condition it seems like time

"\Delta t"

was such that

"a< g"

. This happened because of the drag forces caused by the opened parachute and the influence of air resistance.